OpenStudy (anonymous):

Find all polar coordinates of point P where P = (4,-pi/3)

OpenStudy (anonymous):

@VeritasVosLiberabit plz help

OpenStudy (anonymous):

sure, let me type an explanation for you

OpenStudy (anonymous):

Here are the choices if you need them

OpenStudy (anonymous):

\[(r,\theta)\] is polar coordinate form \[r=\frac{ x }{ \cos (\theta) }=\frac{ y }{ \sin(\theta) }\] \[\theta=\tan ^{-1}(\frac{ y }{ x })\] 

OpenStudy (anonymous):

oh ok what's next?

OpenStudy (anonymous):

plug the coordinates you are given \[(4,-\frac{ \pi }{ 3 })\] and solve for both \[(r,\theta)\] with the equations I've given you

OpenStudy (anonymous):

you have two different options with r which should give you the same answer

OpenStudy (anonymous):

how do i solve for r,theta though? can u show me?

OpenStudy (anonymous):

do you see the equations I wrote above. Use them to solve for both

OpenStudy (anonymous):

ohhhh alright 1 second

OpenStudy (anonymous):

You simply plug in the x and y values and you will have r and theta

OpenStudy (anonymous):

i don't think my answer is right

OpenStudy (anonymous):

r=4sec(theta)=-1/3picsc(theta)

OpenStudy (anonymous):

Tell me what you have for r, I didn't realize the answer also wanted all values of theta

OpenStudy (anonymous):

Ah I think I may have confused you. You can just use 1 equation for r. Either the one with sin or the one with cos

OpenStudy (anonymous):

what value did you get for theta?

OpenStudy (anonymous):

1 second

OpenStudy (anonymous):

hey @ButIneedHelp I'm sorry forget everything I said. The original problem is already in polar

OpenStudy (anonymous):

oh ok so what do we do?

OpenStudy (anonymous):

Let me look back at the problem I didn't realize it was polar coordinates

OpenStudy (anonymous):

Alright thank you

OpenStudy (anonymous):

They basically are looking to see all possible values the line can take so the only thing that you need to find is all the possibilities the angle can be

OpenStudy (anonymous):

Can you show me how to do that?

OpenStudy (anonymous):

sure for one it is best just to look straight at the answers given in this one

OpenStudy (anonymous):

I have a feeling the answer is d, is it?

OpenStudy (anonymous):

we know that the point \[(4,-\frac{ \pi }{ 3 })=(4,-\frac{ \pi }{ 3 }+2n \pi)\]

OpenStudy (anonymous):

No not d

OpenStudy (anonymous):

oh ok damn

OpenStudy (anonymous):

So its either a or c?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Any angle + 2(pi)n will go back to the same position where n is an integer value

OpenStudy (anonymous):

So it's a?

OpenStudy (anonymous):

yes how did you know or is that a guess?

OpenStudy (anonymous):

I guessed to be honest, because the other options have the same beginning to the second parenthesis

OpenStudy (anonymous):

Can you help me with a few more? I'mi almost done

OpenStudy (anonymous):

ok well it is a because if you notice the line has is located where x is -4 now

OpenStudy (anonymous):

it adds 3 pi to get back to the original spot this is hard to explain without a picture

OpenStudy (anonymous):

Make a new thread and I will help you sure